# Stability theory over toroidal or Novikov-type base and canonical modifications

### v2

@@ Abstract.
We set up a generalization of ubiquitous one parameter families in algebraic geometry and its use for stability theories ([Mum65, HL14, AHLH23]) to families over toric varieties, Novikov type rings, or its complex analytic analogues. The language allows us to reformulate degenerations of “irrational” direction in various literatures as canonical objects in a unified manner.
Compatibly, we generalize the stable reduction type theorem for $\Theta$-stratification in [AHLH23] of Langton type to our higher rank setup.
@@ 1. Introduction
In the classical geometric invariant theory [Mum65], the so-called Hilbert-Mumford criterion for stability of an algebro-geometric object, say ${\bf X}$, with respect to an action of a reductive algebraic group $G$, shows that it is enough to analyze degenerations of ${\bf X}$ along actions of $\mathbb{G}_{m}$ which is inside $G$ ($1$-parameter subgroups). The resulted families are isotrivial degenerating families of ${\bf X}$ over $\mathbb{A}^{1}$. In the case when ${\bf X}$ is a polarized projective variety, they are characterized as test configurations by Donaldson [Don02] (cf., also [Don05]).
More precisely speaking, for the Hilbert-Mumford criterion, one associates GIT weight to each $1$-parameter subgroup and their positivity is equivalent to the stability. On the other hand, it is also well-known that we can alternatively think of maximal tori $T$ of $G$, instead of $1$-parameter subgroups, and encode all the data of GIT weights inside $T$ in the weight polytope. The corresponding natural extension of test configurations or the above type isotrivial degenerations, replacing $\mathbb{G}_{m}$ by a maximal torus, is isotrivial family over $T$ and its partial equivariant compactification i.e., toric varieties. At the risk of being misunderstood, a primitive toy model of the perspective in this paper, is such isotrivial degeneration over toric variety of higher dimension.
In this paper, we introduce more extensive algebraic (partly stack-theoretic) generalities with such perspective, which leads to introduce e.g., isotrivial degenerating families over toroidal bases, Novikov type valuation rings of high rank, and their analytic analogues. In particular, focusing on the context of stability and moduli theories of varieties, we introduce the notion of generalized test configurations and their limit versions. Then, together with a companion paper, we show that the generalized notion is useful to treat the following types of degenerations in “irrational direction” in literatures, a more canonical and unified manners:
@@
In our other papers we plan to discuss and explore these one by one, but this paper still focuses on the general framework. In particular, the latter half of this paper gives some generalization of the $\Theta$-stability theory and stable reduction theorem ([HL14] and [AHLH23] respectively), as a core theorem in this paper. Here, recall $\Theta:=[\mathbb{A}^{1}/\mathbb{G}_{m}]$ over a field $k$ is the quotient stack whose coarse moduli is just one point which is used as a source of map to express equivariant isotrivial families over $\mathbb{A}^{1}$. [HL14, AHLH23] introduces the notion of $\Theta$-strata and $\Theta$-stratification on (highly non-separated) algebraic stacks, as a stacky reminicsents of the Hesselink-Kempf-Kirwan-Ness stratification [Kir84, Nes84] and the Harder-Narasimhan filtration of coherent sheaves [HN74]. We call a natural higher rank version of it as higher $\Theta$-strata. Also, as an irrational analogue of DVR, we introduce Novikov type spectra $\overline{\Delta_{\xi}}(R)$ (Definition 2.3). This gives an extension of the valuation ring $R$.
We give a brief outlook of our main theorem but we refer to later Theorem 3.8, 3.15, Corollary 3.10 for the precise statements and generalizations.
Theorem 1.1 (Higher $\Theta$-stable reduction).
Take any quotient stack $\mathcal{M}=[M/G]$ where $M$ is an algebraic $\mathbbm{k}$-scheme and $G$ is a linear algebraic $\mathbbm{k}$-group and let us set $\Delta=\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$ (resp., a complex unit disk $\Delta$, if $\mathbbm{k}=\mathbb{C}$).
@@
Consider any $\mathbbm{k}$-morphism $f\colon\Delta\to\mathcal{M}$ from to $\mathcal{M}$ whose closed point (resp., origin) maps into a closed substack $\mathcal{Z}^{+}$ corresponding higher $\Theta$-strata for the cone $\tau\ni\xi$ (see Definition 3.3) while the generic point (resp., general points) maps outside $\mathcal{Z}^{+}$. We denote the restriction of $f$ to $\mathop{\mathrm{Spec}}\nolimits(K)$ (resp., $\Delta\setminus\{0\}$) as $f^{o}$.
Then, after a finite extension of $R$ (resp., $\Delta$), there is a canonical (for each $\xi$) modification of $f$ as follows:
$f_{{\rm new},\xi}\colon\overline{\Delta_{\xi}}(R)\quad(\text{resp., the universal cover of }\Delta\setminus 0)\to\mathcal{M},$
@@
which still “extends” $f^{o}$ and sends the closed point to a point outside $Z^{+}$, which still isotrivially degenerates to a point in $Z^{+}$.
We call this replacement of $f$ by $f_{{\rm new},\xi}$ as higher $\Theta$-stable reduction and call its limit as higher $\Theta$-stable limit. When $\xi$ is a rational vector, it is nothing but the result of [AHLH23, §6] (“$\Theta$-stable reduction, “$\Theta$-stable limit”). In general case, relations between higher $\Theta$-stable reduction (resp., higher $\Theta$-strata) and the original stable reduction (resp., $\Theta$-stara) are unclear at the moment. Nevertheless, under the assumption of their close connection, we could also apply our method to study certain deformation of $\Theta$-stratification along an algebraic torus and how it affects the $\Theta$-stable limits and show its essentially construtible behaviour (Corollary 3.13).
@@
The framework and theorems in this paper are applied in our companion papers ([Od24a, Od24c]).
@@ 2. Generalities on multi-parameter families
In this section, we generalize some part of the theories of one parameter families and degenerations, which appear in the theory of K-stability or its local analogues for klt singularities (cf., [Li18] and subsequent developments), and the stacky stability theories in [HL14, AHLH23], to those over toroidal base of often higher dimension. Some germs of our theory can be seen in [Spi90, Tei03, DS17, LX18, HL20], for instance. We observe that our language allows re-explanation in a unified manner of those existing K-stability theories.
@@
Further, as we will later explain, the notion allows us to treat the above examples of degenerations (2) and (3) in a (more) canonical manner, to remove the broad ambiguity of choices of $1$-parameter subgroups, except for those with parentheses.
@@ 2.1. Setup of the bases
Since our bases will be often toric geometric, first we set the usual notations for toric geometry.
Notations 1.
We fix a lattice $N\simeq\mathbb{Z}^{r}$ and its rational polyhedral cone $\tau\subset N_{\mathbb{R}}:=N\otimes_{\mathbb{Z}}\mathbb{R}$. The dual lattice of $N$ is denoted as $M:={\rm Hom}(N,\mathbb{Z})$ and we set $M_{\mathbb{R}}:=M\otimes\mathbb{R}$. We denote the algebraic torus $T:=\mathbb{G}_{m}\otimes_{\mathbb{Z}}N$. If we regard $M$ as the character lattice of $T$, we denote by $\chi(\vec{m})$ the character of $T$ which corresponds to each $\vec{m}\in M$. An affine toric partial compactification of $T$ for $\tau$ is denoted by $U_{\tau}(\supset T)$. We write $\tau^{\vee}:=\{x\in M\otimes\mathbb{R}\mid\langle x,\tau\rangle\subset\mathbb{R}_{\geq 0},$ its interior as $\tau^{\vee,o}$, $\mathcal{S}_{\tau}:=\tau^{\vee}\cap M$, and $\mathcal{S}_{\tau}^{o}:=\mathcal{S}_{\tau}\cap\tau^{\vee,o}$.
@@
Under the Notations 1, the global N-ring  [^1] N stands for Novikov as these rings are variants of so-called Novikov ring, in the contexts of Morse theory, Gromov-Witten theory, symplectic geometry, which seem to have different versions in the literatures cf., e.g., [Nov81, Nov82, FOOO09, AJ10, AI21]. over a ring $A$ for $\xi\in N_{\mathbb{R}}$ is the semiring $A[\Gamma_{\geq 0}]$ and we call its spectra $\mathop{\mathrm{Spec}}\nolimits A[\Gamma_{\geq 0}]$ as global N-spectra, which we denote as $\Delta_{\xi}^{\rm gl}(A)$ or simply $\Delta_{\xi}^{\rm gl}$ if the base ring $A$ is obvious from the context. Then, we define the following.
  1. (1)
    Let $I$ be the maximal ideal generated by the monomials corresponding to $\Gamma_{\geq 0}$ and the formal completion of $A[\Gamma_{\geq 0}]$ with respect to $I$ is denoted as $A[[\Gamma_{\geq 0}]]$. With its formal variable $t$, the N-ring over a ring $A$ for $\xi\in N_{\mathbb{R}}$ is its subring defined as
@@
From the perspective of the above proposition, it is naturally to consider the quotient stack $[\Delta_{\xi}^{\rm gl}/T]$ over the category of (non-finite type) schemes, and denote it as $\Delta_{\xi}$ henceforth. For later purpose in the next section, we also introduce a $R$-algebra variant:
Definition 2.3 (N-spectra over $R$).
For $N,\xi$ as notation 1, we fix a local ring $R$ with its non-unit non-zero element $\pi$. We also take and fix $c\in\mathbb{R}_{>0}$, set $\xi^{\prime}:=(\xi,c)\in N_{\mathbb{R}}\times\mathbb{R}_{>0}.$
@@
Under Notationss 1 and Definition 2.1 (ii), take any algebraic $\mathbb{C}$-scheme $X$. Then, any $\mathbb{C}$-morphism (as schemes) $\varphi\colon\Delta_{\xi}^{*,{\rm conv}}(\mathbb{C})\to X$ induces a holomorphic map $\varphi_{\epsilon}^{\rm an}\colon\widetilde{\mathbb{C}}_{\epsilon}^{*}\to X(\mathbb{C})$ for $0<\epsilon\ll 1$.

### v3

@@ Abstract.
We set up a generalization of ubiquitous one-parameter families in algebraic geometry and their use for stability theories ([Mum65, HL14, AHLH23]) to families over toric varieties and their analytic analogues. The language allows us to reformulate degenerations of “irrational” direction in the literature as canonical objects in a unified manner.
Accordingly, we generalize the (semi)stable reduction-type theorem for $\Theta$-stratification in [AHLH23] of Langton type to our higher rank setup. We also establish complex analytic analogue of the results. As an infinitesimal analogue of toric spectrum, we also use Novikov type rings as it gives more canonicity but its use can be avoided logically for readers for readers who prefer not to use such rings.
As applications, we establish the properness part of the moduli of Calabi-Yau cones (cf., also [Od24a]), and also reduce the properness of the moduli of Kähler-Ricci solitons, again to a finite generation type problem in birational geometry.
@@ 1. Introduction
In the classical geometric invariant theory [Mum65], the so-called Hilbert-Mumford criterion for stability of an algebro-geometric object, say ${\bf X}$, with respect to an action of a reductive algebraic group $G$, shows that it is enough to analyze degenerations of ${\bf X}$ along actions of $\mathbb{G}_{m}$ which is inside $G$ ($1$-parameter subgroups). The resulting families are isotrivially degenerating families of ${\bf X}$ over $\mathbb{A}^{1}$. In the case when ${\bf X}$ is a polarized projective variety, they are characterized as test configurations by Donaldson [Don02] (cf., also [Don05]).
More precisely, for the Hilbert-Mumford criterion, one associates a GIT weight to each $1$-parameter subgroup and their positivity is equivalent to the stability. Alternatively, it is well known that one can consider maximal tori $T$ of $G$, instead of $1$-parameter subgroups, and encode all GIT weights data within $T$ in the weight polytope. The corresponding natural extension of test configurations or the above type isotrivial degenerations, replacing $\mathbb{G}_{m}$ by a maximal torus, is isotrivial family over $T$ and its partial equivariant compactification i.e., toric varieties. A simple toy model for the principal objects in this paper is an isotrivial degeneration over a higher-dimensional toric variety.
More generally, we introduce more extensive algebraic (partly stack-theoretic) generalities with such perspective, which leads to introduce e.g., isotrivial degenerating families over toroidal bases, Novikov type valuation rings of high rank, and their analytic analogues. In particular, focusing on the context of stability and moduli theories of varieties, we introduce the notion of generalized test configurations and their limit versions. Then, together with a companion paper, we show that the generalized notion is useful to treat the following types of degenerations in “irrational direction” in literatures, a more canonical and unified manners:
@@
In our §4 and [Od24a, Od24c], we discuss and explore these examples one by one, but until §3 of this paper we focus on the general framework. In particular, our §3 gives some generalization of the $\Theta$-(semi)stability theory and semistable reduction theorem ([HL14] and [AHLH23] respectively), as a core theorem in this paper. Here, recall $\Theta:=[\mathbb{A}^{1}/\mathbb{G}_{m}]$ over a field $k$ is the quotient stack whose coarse moduli is just one point which is used as a source of map to express equivariant isotrivial families over $\mathbb{A}^{1}$. [HL14, AHLH23] introduces the notion of $\Theta$-strata and $\Theta$-stratification on (highly non-separated) algebraic stacks, as a stacky reminicsents of the Hesselink-Kempf-Kirwan-Ness stratification [Kir84, Nes84] and the Harder-Narasimhan filtration of coherent sheaves [HN74]. We call a natural higher $=:r$) version of it as higher $\Theta$-strata by roughly considering degenerations along $\Theta^{r}=[\mathbb{A}^{r}/T]$ with $T:=\mathbb{G}_{m}^{r}$ (temporarily). One can also replace $\mathbb{A}^{r}$ by other affine toric variety $U_{\tau}$ for a cone $\tau$. Also, as an irrational analogue of DVR, we introduce Novikov type spectra $\overline{\Delta_{\xi}}(R)$ (Definition 2.3). This gives an extension of the valuation ring $R$.
To give a feel of our main statements on such higher $\Theta$-stratum, here we show its quick brief outlook, although we refer to later Theorem 3.8, 3.15, Corollary 3.10 for the precise statements and generalizations. Roughly put, it shows the existence of “improvement/modification” of one parameter family, which avoids some “too unstable” locus in a fairly general setup.
For below, we consider a quotient stack $\mathcal{M}=[M/G]$ over $\mathbbm{k}$, where $M$ is an algebraic $\mathbbm{k}$-scheme and $G$ is a linear algebraic $\mathbbm{k}$-group (e.g., ${\rm GL}_{n}$ or algebraic torus). In that case, higher $\Theta$-stratum $\mathcal{Z}^{+}=[Z/G]$ with the closed subscheme $Z\subset M$ and a rational polyhedral cone $\tau\subset N\otimes\mathbb{R}$ roughly means for each point $p\in Z(\mathbbm{k})$, there is its isotrivial degeneration (the generalized Harder-Narasimhan filtration) along the affine toric variety $U_{\tau}\supset T$ with a homomorphism $T\to G$ and $Z^{+}$ is the maximal locus where these structure varies uniformly. Note that our $r$ is not necessarily the rank of $G$.
Theorem 1.1 (Higher $\Theta$-semistable reduction: cf., Corollary 3.10).
Consider a quotient stack $\mathcal{M}$ over a field $\mathbbm{k}$ and its higher $\Theta$-strata $\mathcal{Z}^{+}\subset\mathcal{M}$ for the cone (see Definition 3.3).
@@
Take any $\mathbbm{k}$-morphism $f\colon\Delta=\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]\to\mathcal{M}$ such that its closed point maps into $\mathcal{Z}^{+}\subset\mathcal{M}$ while the generic point maps outside $\mathcal{Z}^{+}$. We denote the restriction of $f$ to $\mathop{\mathrm{Spec}}\nolimits(K)$ as $f^{o}$.
Then, for a finite extension $R^{\prime}(\simeq\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t^{1/d}]]$ for some $d\in\mathbb{Z}_{>0}$) of $R$, there exists a modification of $f$ as follows:
$f_{{\rm new}}\colon\mathop{\mathrm{Spec}}\nolimits(R^{\prime})\to\mathcal{M},$
@@
which still extends $f^{o}$ but now sends the closed point $c_{R^{\prime}}$ to a point outside $Z^{+}$.
If there are finite strata $\{\mathcal{Z}_{i}^{+}\}_{i}$ inside $\mathcal{M}$ which form what we call higher $\Theta$-stratification (Definition 3.17), extending the case #$\#\{i\}=1$ above, we can iterativly apply the above modification (see Theorem 3.18) to show the properness of the complement of $\mathcal{Z}_{i}^{+}$s, which often admit moduli interpretations (see Appendix 4 for specific examples). Indeed, in the moduli contexts, $\mathcal{Z}_{i}$s often appear as unstable locus in the spirit of [Mum65] so that their complement is nothing but the locus which parametrizes semistable objects.
Here are some technical comments in order. For Theorem 1.1 mentioned above, we also refine the statements in the main body of this text from at least the following four different aspects.
The first aspect is about the new base $R^{\prime}$. There are many choices of the pair of DVR $R^{\prime}$ and $f_{\rm new}$ as a family one can take in our construction, though $f_{\rm new}(c_{R^{\prime}})$ is more canonical. Indeed, they roughly corresponds to choice of integral points in a certain rational polyhedral cone so that there is no canonical choice and there are infinite possibilities. Nevertheless, if we fix a possibly irrational vector $\xi(\neq 0)\in\tau$ and allow $R^{\prime}$ to be non-discrete valuation ring in general, there is more canonical choice of $R^{\prime}$ as a bigger valuation ring for each $\xi$ and we expect its importance for future study. That is the Novikov type ring $\overline{\Delta_{\xi}}(R)$ we mentioned above, as we introduce and discuss in §2. We also discuss various variants of related Novikov type rings and relation with toric geometry, motivated by the above reason.
Secondly, we also prove its complex analytic version (Theorem 3.15) by using the universal covering of punctured disk or real ray in the unit disk (cf., Definition 2.4). We call this replacement of $f$ as higher $\Theta$-semistable reduction and call its limit as higher $\Theta$-semistable limit. When $\xi$ is an integral vector, our result partially recovers [AHLH23, §6, Theorem 6.3, Theorem 6.5] by more scheme theoretic approach.
Thirdly, in the case when $\mathcal{M}$ is of the form $[M/T]$, we prove certain uniqueness of the above modification in a weak sense. See (iii) of Theorem 3.8 for the details.  [^1] We hope this will be benefitial for further study of bubbling theory [Od24c].
Lastly, but almost obviously, if there are certain stratifications of $\mathcal{M}$ consists of higher $\Theta$-strata, we can simply repeat the above modification to give a stronger result (Theorem 3.18) as it was in [AHLH23].
@@
The framework and theorems in this paper are applied in our §4 and our companion papers ([Od24a, Od24c]), which are mostly in the context of Kähler geometry and algebraic theory related to K-stability.
Notes added in version 3
About $8$ and a half month after the first version of this paper, we saw some developments i.e., [BHLINK25] appeared, but with a different motivation. At the end of the paper, put in our version 3 on arXiv, we discuss the overlap, interesting relations and certain differences between this paper and [BHLINK25].
@@ 2. Generalities on multi-parameter families
This section is of preparatory nature. We generalize some part of the theories of one parameter families and degenerations, which appear in the theory of K-stability or its local analogues for klt singularities (cf., [Li18] and subsequent developments), and the stacky stability theories in [HL14, AHLH23], to those over toroidal base of often higher dimension. Some germs of our theory can be seen in [Spi90, Tei03, DS17, LX18, HL20], for instance. We observe that our language allows re-explanation in a unified manner of those existing K-stability theories.
@@
Further, as we will later explain, the notion allows us to treat the above examples of degenerations ($\mathrm{ii}$) and ($\mathrm{iii}$) in a (more) canonical manner, to remove the broad ambiguity of choices of $1$-parameter subgroups, except for those with parentheses.
@@ 2.1. Setup of the bases
Since our bases will be often toric geometric, first we set the usual notation for toric geometry.
Notation 1.
We fix a free abelian group (lattice) $N$ of rank $r$ and its rational polyhedral cone $\tau\subset N_{\mathbb{R}}:=N\otimes_{\mathbb{Z}}\mathbb{R}$. The dual lattice of $N$ is denoted as $M:={\rm Hom}(N,\mathbb{Z})$ and we set $M_{\mathbb{R}}:=M\otimes\mathbb{R}$. We define the algebraic split torus $T=\mathbb{G}_{m}\otimes_{\mathbb{Z}}N$ as ${\rm Spec}\mathbb{C}[M]$, where $\mathbb{C}[M]$ means the group ring. If we regard $M$ as the character lattice of $T$, we denote by $\chi(\vec{m})$ the character of $T$ which corresponds to each $\vec{m}\in M$. An affine toric partial compactification of $T$ for $\tau$ is denoted by $U_{\tau}(\supset T)$. We write $\tau^{\vee}:=\{x\in M\otimes\mathbb{R}\mid\langle x,\tau\rangle\subset\mathbb{R}_{\geq 0},$ its interior as $\tau^{\vee,o}$, $\mathcal{S}_{\tau}:=\tau^{\vee}\cap M$, and $\mathcal{S}_{\tau}^{o}:=\mathcal{S}_{\tau}\cap\tau^{\vee,o}$.
@@
Under the Notation 1, the global N-ring  [^2] N stands for Novikov as these rings are variants of so-called Novikov ring, in the contexts of Morse theory, Gromov-Witten theory, symplectic geometry, which seem to have different versions in the literatures cf., e.g., [Nov81, Nov82, FOOO09, AJ10, AI21]. over a ring $A$ for $\xi\in N_{\mathbb{R}}$ is the semiring $A[\Gamma_{\geq 0}]$ and we call its spectra $\mathop{\mathrm{Spec}}\nolimits A[\Gamma_{\geq 0}]$ as global N-spectra, which we denote as $\Delta_{\xi}^{\rm gl}(A)$ or simply $\Delta_{\xi}^{\rm gl}$ if the base ring $A$ is obvious from the context. Then, we define the following.
  1. ($\mathrm{i}$)
    Let $I$ be the ideal of $A[\Gamma_{\geq 0}]$ generated by the monomials corresponding to $\Gamma_{\geq 0}$ and the formal completion of $A[\Gamma_{\geq 0}]$ with respect to $I$ is denoted as $A[[\Gamma_{\geq 0}]]$. With its formal variable $t$, the N-ring over a ring $A$ for $\xi\in N_{\mathbb{R}}$ is a subring of $A[[\Gamma_{\geq 0}]]$ defined as
@@
From the perspective of the above proposition, it is naturally to consider the quotient stack $[\Delta_{\xi}^{\rm gl}/T]$ over the category of (non-finite type) schemes, and denote it as $\Delta_{\xi}$ henceforth. For later purpose in the next section, we also introduce a $R$-algebra variant. More specifically, it will appear as a natural (canonical) base of a modification of morphism from ${\rm Spec}(R)$ to an algebraic stack, in our higher analogue of semistable reduction type theorem (Theorem 3.8, Corollary 3.10).
Definition 2.3 (Novikov-type spectra over $R$).
For $N,\xi$ as notation 1, we fix a local ring $R$ and also fix its non-unit non-zero element $\pi$. We also take and fix $c\in\mathbb{R}_{>0}$, set $\xi^{\prime}:=(\xi,c)\in N_{\mathbb{R}}\times\mathbb{R}_{>0}.$
@@
which is dual to $N^{\prime}\hookrightarrow N\oplus\mathbb{Z}$. Then, consider the affine scheme over $R$ as
(1) $U_{\tau^{\prime}}(R):=\mathop{\mathrm{Spec}}\nolimits(R[M^{\prime}\cap\tau^{\prime\vee}]/(\pi-\chi^{p(\vec{0},1)})),$
and its projective limit
@@
Morally speaking, it is global in the direction of $\mathbb{G}_{m}\otimes N^{\prime}$-action but it is only local in $R$ direction, which is the reason why we avoid the superscript gl in this case and just use the overline. Here, $\tau^{\prime\vee}$ means the dual cone of $\tau^{\prime}$. As $p(\vec{0},1)\in M^{\prime}\cap\tau^{\prime\vee},$ $\chi^{p(\vec{0},1)}$ means the corresponding monomial (character) of $R[M^{\prime}\cap\tau^{\prime\vee}]$.
We also note that the above ring $R[M^{\prime}\cap\tau^{\prime\vee}]/(\pi-\chi^{p(\vec{0},1)})$ is a quotient ring of $(R[X])[\tau^{\vee}\cap M]/(X\chi^{p(\vec{0},1)}-\pi)$, where $X$ is an extra variable and the quotient is by the ideal $(X-1)$. We also define
@@
These and their further extensions will appear in the proofs of semistable reduction type theorems in §3.2 and later. See Definition 3.5, Remark 3.7 Theorem 3.15, and Corollary 3.10.
@@
Under Notation 1 and Definition 2.1 (ii), take any algebraic $\mathbb{C}$-scheme $X$. Then, any $\mathbb{C}$-morphism (as schemes) $\varphi\colon\Delta_{\xi}^{*,{\rm conv}}(\mathbb{C})\to X$ induces a holomorphic map $\varphi_{\epsilon}^{\rm an}\colon\widetilde{\mathbb{C}}_{\epsilon}^{*}\to X(\mathbb{C})$ for $0<\epsilon\ll 1$.